OpenStudy (anonymous):

What volume in liters of Cl2O gas at STP can be produced from 422 cm3 of chlorine gas at STP and 9.65 g of HgO?

OpenStudy (anonymous):

2 Cl2 + HgO → Cl2O + HgCl2 (0.422 L Cl2) / (22.414 L/mol) = 0.018828 mol Cl2 (9.65 g HgO) / (216.5894 g HgO/mol) = 0.044554 mol HgO 0.018828 mole of Cl2 would react completely with 0.018828 x (1/2) = 0.009414 mole of HgO, but there is more HgO present than that, so HgO is in excess and Cl2 is the limiting reactant. (0.018828 mol Cl2) x (1 mol Cl2O / 2 mol Cl2) x (22.414 L/mol) = 0.211 L Cl2O